What capacitor in series with a 100 Ω resistor and a 19 mH inductor will give a resonance frequency of 1000 Hz ? Express your answer in microfarads.
Answer #1

Resonance frequency, f = 1/ 2*pi*sqrt(L*C)

sqaring both sides,

f2=1/4 *pi*pi*L*C

C=1/4 *pi*pi*L*f2

C=1/4*3.14*3.14*19*10-3*1000*1000

C= 1.33*10-6 Farad

C=1.33 micro farad